OpenStudy (dumbcow):

Given a sphere perfectly inscribed inside of a regular square pyramid. Find the ratio of the volume of sphere to volume of pyramid. good luck (exact answers only)

OpenStudy (dumbcow):

drawing of base and triangular face 

OpenStudy (raden):

pi (3sqrt(3)-5) : 2 ??

OpenStudy (dumbcow):

congrats! not in same form as my solution, but numerically the same \[ratio = \frac{\sqrt{2}}{16}(\sqrt{6}-\sqrt{2})^{3} \pi : 1\]

OpenStudy (raden):

yesss !! :)